OpenStudy (anonymous):

Evaluate: 8 + g ÷ h • 4, when g = 9 and h = 3.

OpenStudy (anonymous):

17/12

OpenStudy (anonymous):

be careful to use parenthesis, the expression is ambiguous

OpenStudy (sriram):

=8+3*4 =8+12 =20